Find the least common denominator (LCD) for fractions with the given denominators.
1. 14 and 21
2. 3, 4, and 5

Answer by mananth(16946)

1. 14 and 21
14= 2*7
21=3*7
common factor = 7
LCD = common factor * non common factors
LCD = 7 * 3*2
LCD= 42
...
2. 3, 4, and 5
60
...
m.ananth@hotmail.ca